OpenStudy (anonymous):

A motorized scooter has tires with a radius of 3 1/2 inches.vThe tires make one revolution every 1/10 second. Find the speed of the scooter in miles per hour. (hint: THe speed of an object spinning in a circle is equal to the circumference divided by the time it takes to complete one revolution)

jigglypuff314 (jigglypuff314):

Hello @stylestar5678 and Welcome to OpenStudy :) first find the circumference of the tire cricumference = 2*pi*r = ?

jigglypuff314 (jigglypuff314):

then know that there are 63360 inches in one mile and that there are 3600 seconds in one hour

OpenStudy (anonymous):

im a bit confused on how to solve it and put it all together

jigglypuff314 (jigglypuff314):

that's ok, first find the circumference can you do that? :)

OpenStudy (anonymous):

how?

jigglypuff314 (jigglypuff314):

circumference = 2*pi*r = 2*(pi)*(3.5inches) = ?

OpenStudy (anonymous):

so its about 21.98?

OpenStudy (anonymous):

is it?

jigglypuff314 (jigglypuff314):

21.99 yes :)

OpenStudy (anonymous):

now what?

jigglypuff314 (jigglypuff314):

so now you have the tires spinning at the speed of 21.99 inches ------------ 0.1 secs do you see where I got those numbers? :)

OpenStudy (anonymous):

yes

jigglypuff314 (jigglypuff314):

now we want to get those inches to become miles and those seconds to become hours

OpenStudy (anonymous):

proportions?

jigglypuff314 (jigglypuff314):

21.99 inches 1 mile 3600 secs ------------ x ------------ x --------- 0.1 secs 63360 inches 1 hour this would get the seconds to cancel out and leave hours and the inches to cancel out and leave miles :)

OpenStudy (anonymous):

so I just multiply them?

jigglypuff314 (jigglypuff314):

yep :) tell me what you get ^_^

OpenStudy (anonymous):

0ok im doing it now

OpenStudy (anonymous):

ok I got 10764 ------- 6336

OpenStudy (anonymous):

is it right?

jigglypuff314 (jigglypuff314):

hmmm I think 21.99 times 3600 = 79164 ?

OpenStudy (anonymous):

your right i just checked on the calculator. I guess I cant rely on bings calculator anymre..what do I do now?

jigglypuff314 (jigglypuff314):

well 79164 divided by 6336 = the answer you need :)

OpenStudy (anonymous):

12.49?

jigglypuff314 (jigglypuff314):

yes! that's what I got! :D

OpenStudy (anonymous):

so the answer is 12.49 miles per hour?

jigglypuff314 (jigglypuff314):

correct! ^_^

OpenStudy (anonymous):

thanks you so much!

jigglypuff314 (jigglypuff314):

hehe you're welcome :)

OpenStudy (anonymous):

:) bye

OpenStudy (anonymous):

are you sure thts the answer, my friend got 0.063

jigglypuff314 (jigglypuff314):

it should be right

OpenStudy (anonymous):

kk thanks

jigglypuff314 (jigglypuff314):

do you have the textbook in front of you? if it is an odd number problem, the answer should be in the back of the book